OpenStudy (anonymous):

to solve by completing the square, what value should you add to each side of the equation? x^2+12x=-3

OpenStudy (hyper):

6^2 = 36 should be added to both sides to balance the equation

OpenStudy (anonymous):

huh?

OpenStudy (hyper):

(x^2 +12x) = (x + 6)^2 -36 Hence, both sides have to be added with 36 so that they balance: From there, this becomes : (x^2 + 12x) + 36 = -3 +(36) and so x^2 +12x + 36 is just really (x + 6)^2

OpenStudy (anonymous):

i still dont get it. :(

OpenStudy (anonymous):

sorry, I'm stressed I have been doing this all week. I'm fried. :(

OpenStudy (hyper):

ok, for the left hand side, to solve this quadratic equation, it has to be made into a square which is why this process is called completing the square. To make x^2 + 12x into a square, it becomes (x + 12/2)^2 = (x + 6) ^2. However as the left side now has an extra 36 than it had before, 36 or 6^2 has to be also added to the right hand side (-3) to balance the equation.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

12/2 = 6 6^2 = 36 So, you add 36 x^2+12x+36=-3+36 (x+6)^2 = 39

OpenStudy (anonymous):

@Hyper

OpenStudy (anonymous):

36** not 39